# Think How to Think: Mitigating Overthinking with Autonomous Difficulty Cognition in Large Reasoning Models

### v2

@@ Introduction
We extend difficulty-hypnosis of Stage 1 to redundancy-hypnosis, guiding to identify redundant structures within the reasoning process on hard tasks.
@@
where is the difficulty-hypnosis introduced in Stage 1, are valid reasoning chunks to be kept, is the unnecessary reflection structure inside the reasoning process without effective contribution to the reasoning process. These chunks often carry out backtracking, self-verification, or exploration, but unfortunately, are nothing more than superfluous overhead. is a similar reflection structure; nonetheless, it primarily manifests in the final phase of the reasoning process, continuously iterating until attains the maximum output length.
@@
wherein the redundancy hypnosis is defined as follows: “hypnosis Everything seems ok, let’s move on./hypnosis”. We use GPT-4 (OpenAI et al. 2024b) to implement the determinator.
@@
In case of , on one hand, the training set lacks a sufficient number of loop structures; on the other hand, excessively deferred looping hinders the generation of shorter-length samples. Thus, we select intermediate chunk with words like “Wait”, “However”, “Alternatively” and iterate it as a simulation of looping, followed by hypnosis injection and truncation:
@@
wherein the looping hypnosis is defined as follows: “hypnosis Oh, I’m stuck in a loop. Time to break out. /hypnosis”. The choice of iterate number can be random, and we also try a probe-like approach (Zhang et al. 2025). We use DeepSeek-R1 to identify chunks before which the prior content can already induce the correct answer, referred to as completeness requirements. The last-layer hidden states of the selected chunk tail is extracted to train a simple two-layer MLP as Determinator. Experiments indicate it yields similar results to the random approach, so we prefer the latter.
Notably, we adopt truncation instead of mere substitution, as truncation allows for the construction of shorter samples. Meanwhile, generation continuity is ensured by remaining untruncated samples in the dataset, since we only apply Stage 2 to part of the samples in . Stage 2 execution is outlined in Fig. 4.
Experiments
Experiment Setup
Models.
Dataset Construction.
@@
For dataset construction, we mainly adopt mathematical reasoning tasks, which are friendly for difficulty quantization. In stage 1, the GSM8K and MATH datasets are combined after filtering out errors and injecting difficulty hypnosis. In stage 2, redundancy and looping hypnosis are injected into difficult samples. The composition ratio of this heterogeneous dataset is 2:1:1.
Baselines.
We compare TH2T with two prompt-based methods, i.e., D-Prompt and NoThinking (Ma et al. 2025a), and two fine-tuning methods, i.e., Tokenskip (Xia et al. 2025), CoT-Valve (Ma et al. 2025b). D-Prompt directly promotes LRM to regulate its output length based on self-judgment of task difficulty, through difficulty reminder in the prompt, as discussed in Preliminary Part. The remaining methods are from recent publications.
| Methods | GSM8K | MATH-500 | ||||||||||
| Acc | Gain | Len | Reduc | Latency | SpdUp | Acc | Gain | Len | Reduc | Latency | SpdUp | |
| DeepSeek-R1-Distill-Qwen-7B | ||||||||||||
| Original | 91.51 | 1057 | 20.29 | 85.2 | 2857 | 87.69 | ||||||
| [1pt/2pt] D-Prompt | 90.59 | 0.92 | 906 | 14.3% | 16.38 | 1.24 | 86.0 | 0.8 | 2726 | 4.6% | 81.54 | 1.08 |
| NoThinking | 88.47 | 3.04 | 243 | 76.9% | 3.34 | 6.07 | 79.4 | 5.8 | 699 | 75.5% | 11.81 | 6.43 |
| [1pt/2pt] TokenSkip | 88.01 | 3.50 | 638 | 39.6% | 10.64 | 1.91 | 80.2 | 5.0 | 2109 | 26.2% | 53.94 | 1.62 |
| CoT-Valve | 87.90 | 3.61 | 235 | 77.8% | 3.31 | 6.13 | 78.6 | 6.6 | 747 | 73.9% | 12.31 | 7.12 |
| (w/o D.H.) | 90.83 | 0.68 | 503 | 47.4% | 7.80 | 2.60 | 85.2 | 0.0 | 1906 | 33.2% | 46.68 | 1.88 |
| (w/o R.H.) | 91.96 | 0.48 | 397 | 62.4% | 5.99 | 3.39 | 86.0 | 0.8 | 2265 | 20.7% | 60.22 | 1.46 |
| TH2T | 92.87 | 1.36 | 275 | 74.0% | 3.88 | 5.23 | 86.8 | 1.6 | 1772 | 38.0% | 41.65 | 2.11 |
| DeepSeek-R1-Distill-Qwen-14B | ||||||||||||
| Original | 93.63 | 584 | 9.38 | 86.6 | 2306 | 62.65 | ||||||
| [1pt/2pt] D-Prompt | 93.91 | 0.28 | 606 | 3.9% | 9.95 | 0.94 | 86.6 | 0.0 | 2297 | 0.3% | 62.60 | 1.01 |
| NoThinking | 88.41 | 3.48 | 206 | 64.9% | 3.06 | 3.07 | 76.0 | 10.6 | 612 | 73.5% | 10.05 | 6.23 |
| [1pt/2pt] TokenSkip | 89.30 | 4.33 | 314 | 46.2% | 4.79 | 1.96 | 73.2 | 13.4 | 1356 | 41.2% | 28.11 | 2.23 |
| CoT-Valve | 92.70 | 0.90 | 526 | 9.9% | 8.13 | 1.15 | 84.2 | 2.4 | 2111 | 8.4% | 53.95 | 1.16 |
| (w/o D.H.) | 93.69 | 0.06 | 338 | 42.2% | 4.82 | 1.95 | 86.2 | 0.4 | 1549 | 32.8% | 33.93 | 1.85 |
| (w/o R.H.) | 93.47 | 0.16 | 299 | 51.2% | 4.34 | 2.16 | 85.8 | 0.8 | 1706 | 26.0% | 39.83 | 1.57 |
| TH2T | 94.08 | 0.45 | 246 | 58.1% | 3.38 | 2.78 | 86.4 | 0.2 | 1343 | 41.8% | 27.84 | 2.25 |
| DeepSeek-R1-Distill-Qwen-32B | ||||||||||||
| Original | 95.60 | 717 | 12.14 | 87.2 | 2357 | 63.02 | ||||||
| [1pt/2pt] D-Prompt | 95.60 | 0.00 | 663 | 7.6% | 11.12 | 1.09 | 86.8 | 0.4 | 2198 | 6.8% | 57.11 | 1.10 |
| NoThinking | 93.86 | 1.74 | 228 | 68.2% | 3.25 | 3.74 | 80.2 | 7.0 | 659 | 72.0% | 10.98 | 5.74 |
| [1pt/2pt] TokenSkip | 96.54 | 0.94 | 423 | 41.0% | 6.28 | 1.93 | 79.8 | 7.4 | 1567 | 33.5% | 34.26 | 1.84 |
| CoT-Valve | 96.10 | 0.50 | 781 | 8.9% | 12.56 | 0.97 | 85.3 | 1.9 | 2263 | 3.9% | 60.21 | 1.05 |
| (w/o D.H.) | 96.08 | 0.48 | 361 | 49.7% | 5.47 | 2.22 | 87.4 | 0.2 | 1863 | 20.9% | 45.18 | 1.39 |
| (w/o R.H.) | 95.75 | 0.15 | 323 | 55.0% | 4.84 | 2.51 | 86.8 | 0.4 | 2114 | 10.3% | 53.99 | 1.17 |
| TH2T | 97.04 | 1.44 | 263 | 63.3% | 3.51 | 3.46 | 88.0 | 0.8 | 1733 | 26.5% | 40.13 | 1.57 |
| Bench | Difficulty Level | Cognition | |
| Original | TH2T | ||
| GSM8K | all easy | 14.9% | 98.3% |
| MATH | lv.1-2 (easy) | 9.8% | 93.4% |
| lv.4-5 (hard) | 12.0% | 91.5% | |
| AIME2024 | all hard | 16.7% | 96.7% |
| OmniMath | lv.1-2 (easy) | 20.0% | 96.7% |
| lv.4-5 (hard) | 13.3% | 93.3% | |
Evaluations.
We evaluate on 4 mathematical benchmarks (GSM8K, MATH, AIME2024 and Omni-MATH) and 1 QA benchmark GPQA (Rein et al. 2024). Metrics include accuracy, average number of output tokens, and inference latency per sample. We also report the success rate of the difficulty awareness, redundancy, and looping optimization.
The detailed experiment settings are in the appendix.
Experimental Results
Main Results.
@@
The performance of TH2T on GSM8K and MATH is illustrated in Tab. 1. Key observations include: 1) TH2T exhibits minimal accuracy degradation; more precisely, most settings witness slight improvement. 2) TH2T achieves the best reduction in reasoning length among methods with stable accuracy—up to 74% on GSM8K and 38% on MATH with the 7B model, translating to over 5 and 2 latency speedups, respectively. 3) Greater gains on easier tasks, validating TH2T’s ability to adapt reasoning effort based on task difficulty. 4) These observations are consistent across 7B, 14B and 32B models. As shown in Fig. 1, other methods (e.g., NoThinking and CoT-Value) achieve comparable compression ratios in token length but at the cost of severe accuracy degradation (often 10 points), underscoring TH2T’s robustness in balancing efficiency and reliability.
Difficulty Cognition.
Does TH2T present better difficulty cognition?. As shown in Tab. 2, the entirety of GSM8K, MATH lv.1-2, and Omni-Math lv.1-2 are set as “Easy” while the remaining are set as “Hard”. The original R1-Distilled-7B model is prompted to predict its own understanding of problem difficulty and choose between “Easy”, “Medium”, and “Hard”. TH2T can response with native self-hypnosis in the prefix like “This is a simple question” or “It seems difficult”, which indicates the cognition of task difficulty as “Easy” or “Hard”. As shown in Tab. 2 and Fig. 2, the original LRM demonstrates ambiguous recognition of task difficulty, uniformly categorizing tasks of varying complexity as medium level. Whereas, our approach exhibits a clear distinction with high precision over 90+%, validating the effectiveness of the first stage of TH2T.
Redundancy Cognition.
@@
Does TH2T present better redundancy cognition? Tab. 3 shows the average number of reflective and looping structures. Comparison is made between R1-Distilled-7B and Ours-7B. We observe that: 1) TH2T effectively suppresses the number of reflection chunks by 10 on easy tasks and 3 on hard tasks, suggesting that extensive reflection is unnecessary for strong reasoning. 2) The looping problem in TH2T undergoes substantial restraint. Notably, 59.5% of incorrect answers by the original model on MATH exhibit tail recursion, which is mitigated to 21.4% in ours. As shown in Fig. 6(b), we observe zero loop structure in GSM8K. This distinct pattern of short-length concentration underscores the efficacy of Stage-2 of TH2T, presenting a more concise reasoning structure.
@@
To validate the generalization on OOD tasks, we evaluate on 2 more mathematical benchmarks and 1 QA benchmark. As shown in Fig. 5, accuracy maintenance and token length reduction are consistent.
Results of Llama series model are in the appendix.
Ablation Study
Ablation on Stages.
@@
We explore two more settings: 1) without Difficulty Hypnosis in Stage-1, and 2) without Redundancy Hypnosis in Stage-2, which are annotated as w/o D.H. and w/o R.H. in Tab. 1. From the perspective of response reduction, we observe that D.H. introduces more gain on easy task GSM8K than on hard task MATH (e.g., 26.6% vs 4.8%, on 7B model), while R.H. exerts a greater impact on MATH (11.6% vs 17.3%). As we illustrate during method formulation, we design Difficulty Cognition Injection Stage to build native perception on easy tasks and Redundancy Cognition Injection Stage to mitigate non-essential structures that frequently occur in hard tasks. These results demonstrate the rationality of our two-stage strategy.
Ablation on Starting Model.
Instead of training on LRMs, we also try directly starting from a Short-CoT model with the same data in Tab. 4, which reveals a notable failure in performance and response length. This motivates us to include an underlying Stage-0 in Fig. 4 to start from an LRM, instead of teaching a Short-CoT model from scratch.
| Method | GSM8K | MATH-500 | ||
| Accuracy | Length | Accuracy | Length | |
| Qwen2.5-7B | 86.20 | 280 | 62.80 | 629 |
| [1pt/2pt] TH2T | 72.48 | 238 | 66.20 | 2333 |
| R1-Distill-7B | 91.51 | 1057 | 85.2 | 2857 |
| [1pt/2pt] TH2T | 92.87 | 275 | 86.8 | 1772 |
Ablation on Training Steps.
We examine the dynamics of accuracy and average token length during the training process (figure in appendix), and obtain two interesting findings: 1) As training steps increase, accuracy initially declines, then returns to the original level, with occasional positive gains during the process. 2) For hard tasks like MATH, the response length exhibits a reverse increase after exceeding a certain number of steps. This inspires us to set the training steps to 6400 to prevent model degradation.
Conclusion

### v3

@@ 4.2 Stage 2: Redundancy Cognition Injection
We then extend difficulty-hypnosis to redundancy-hypnosis, guiding to identify redundant structures within the reasoning process on hard tasks.
@@
where is the difficulty-hypnosis introduced in Stage 1, are valid reasoning chunks to be kept. is the superfluous reflection structure inside the reasoning process without effective contribution to the reasoning process. is a similar reflection structure; nonetheless, it primarily manifests in the final phase of the reasoning, and the model enters an unproductive loop (trap), relentlessly iterating until reaches the maximum output length. This practice severely degrades inference efficiency, as latency often scales super-linearly with sequence length, and the vast majority of problems do not require such exhaustive thought.
@@
wherein the Redundancy Hypnosis is defined as follows: “hypnosis Everything seems ok, let’s move on./hypnosis”. We use GPT-4 OpenAI et al. (2024b) to implement the determinator. In case of , the training set lacks a sufficient number of loop structures; meanwhile, excessively deferred looping hinders the generation of shorter-length samples. Thus, we select intermediate chunk with words like “Wait”, “However”, “Alternatively” and iterate it as a simulation of looping, followed by hypnosis injection and truncation:
@@
wherein the looping hypnosis is defined as follows: “hypnosis Oh, I’m stuck in a loop. Time to break out. /hypnosis”. To determine the optimal timing for truncation, we leverage DeepSeek-R1 to justify the earliest point—a condition we term the completeness requirement—where the preceding content becomes sufficient to derive the correct answer. The choice of iterate number can be random, and we also try a probe-like approach Zhang et al. (2025). The last-layer hidden states of the selected chunk tail are extracted to train a simple two-layer MLP as Determinator. Experiments indicate it yields similar results to the random approach, so we prefer the latter. Notably, we adopt truncation instead of mere substitution, as truncation allows for the construction of shorter samples. Meanwhile, generation continuity is ensured by remaining untruncated samples in the dataset, since we only apply Stage 2 to part of the samples in .
5 Experiments
5.1 Experiment Setup
Models. We do experiments on Long-CoT models, including DeepSeek-R1-Distilled-Qwen-2.5-7B/14B/32B and -Llama-8B Guo et al. (2025). During data construction, the adopted Short-CoT model are Qwen2.5-7B/14B/32B Qwen et al. (2025) and Llama-3.1-8B-Instruct Grattafiori et al. (2024). They are and , respectively.
@@ 5.1 Experiment Setup
Datasets. For dataset construction, we mainly adopt mathematical reasoning tasks, which are friendly for difficulty quantization. In stage 1, the GSM8K and MATH datasets are combined after filtering out errors and injecting difficulty hypnosis. In stage 2, redundancy and looping hypnosis are injected into difficult samples. The composition ratio of this heterogeneous dataset is 2:1:1, yielding samples. We evaluate on 4 mathematical tasks (GSM8K, MATH, AIME2024 and Omni-MATH) and 1 QA benchmark GPQA Rein et al. (2024). Metrics include accuracy, average number of output tokens, and inference latency per sample. We also report the success rate of the difficulty awareness, redundancy, and looping optimization.
Baselines. We compare TH2T with the following methods: (1) D-Prompt (prompt-based) directly regulates response length based on self-judgment of task difficulty via the Difficulty Reminder in Prompt (in Sec.3.2). (2) NoThinking Ma et al. (2025a) (output-based) bypasses the thinking process via prompting for all difficulty-level questions. (3) Tokenskip Xia et al. (2025) (model-based) fine-tunes the model to prune non-essential tokens during inference, regardless of task difficulty. (4) CoT-Valve Ma et al. (2025b) (model-based) learns a controllable direction in the parameter space to steer generation length based on task difficulty. (5) AdaCtrl111We re-implemented all baselines except AdaCtrl (citing its original 7B results due to reproduction complexity). Huang et al. (2025) (model-based) uses reinforcement learning and allocates reasoning budgets based on self-assessed problem difficulty, which is more of a label-driven distillation.
| Methods | GSM8K | MATH-500 | AIME2024 | OmniMath | GPQA | ||||||||||
| Acc. | Len. | Reduc. | Acc. | Len. | Reduc. | Acc. | Len. | Reduc. | Acc. | Len. | Reduc. | Acc. | Len. | Reduc. | |
| DeepSeek-R1-Distill-Qwen-7B | |||||||||||||||
| Original | 91.51 | 1057 | 85.2 | 2857 | 50.0 | 10570 | 45.0 | 5736 | 32.8 | 5349 | |||||
| [1pt/2pt] D-Prompt | 90.59 | 906 | 14.3% | 86.0 | 2726 | 4.6% | 46.7 | 10209 | 3.4% | 45.0 | 5398 | 5.9% | 30.3 | 5117 | 4.3% |
| NoThinking | 88.47 | 243 | 76.9% | 79.4 | 699 | 75.5% | 40.0 | 4134 | 60.9% | 40.0 | 2083 | 63.7% | 29.3 | 2084 | 61.0% |
| TokenSkip | 88.01 | 638 | 39.6% | 80.2 | 2109 | 26.2% | 40.0 | 5559 | 47.4% | 38.3 | 3461 | 39.7% | 28.8 | 4001 | 25.2% |
| CoT-Valve | 87.90 | 235 | 77.8% | 78.6 | 747 | 73.9% | 43.3 | 5871 | 44.5% | 41.7 | 3311 | 42.3% | 28.3 | 3226 | 39.7% |
| AdaCtrl | 90.98 | 349 | - | 74.0 | 3196 | - | 21.3 | 16889 | - | - | - | - | - | - | - |
| [1pt/2pt] (w/o D.H.) | 90.83 | 503 | 47.4% | 85.2 | 1906 | 33.2% | 50.0 | 8213 | 22.3% | 46.7 | 3571 | 37.7% | 33.3 | 4372 | 18.3% |
| (w/o R.H.) | 91.96 | 397 | 62.4% | 86.0 | 2265 | 20.7% | 46.7 | 8684 | 17.8% | 48.3 | 4247 | 26.0% | 32.8 | 4481 | 16.2% |
| TH2T | 92.87 | 275 | 74.0% | 86.8 | 1772 | 38.0% | 50.0 | 7490 | 29.1% | 48.3 | 3377 | 41.1% | 32.8 | 4166 | 22.1% |
| DeepSeek-R1-Distill-Qwen-14B | |||||||||||||||
| Original | 93.63 | 584 | 86.6 | 2306 | 50.0 | 9900 | 46.7 | 5373 | 50.0 | 3847 | |||||
| [1pt/2pt] D-Prompt | 93.91 | 606 | 3.9% | 86.6 | 2297 | 0.3% | 50.0 | 10408 | 5.1% | 41.7 | 4994 | 7.1% | 49.0 | 3946 | 2.6% |
| NoThinking | 88.41 | 206 | 64.9% | 76.0 | 612 | 73.5% | 43.3 | 3735 | 62.3% | 43.3 | 1944 | 63.8% | 45.5 | 1454 | 62.2% |
| TokenSkip | 89.30 | 314 | 46.2% | 73.2 | 1356 | 41.2% | 46.7 | 6772 | 31.6% | 40.0 | 3158 | 41.2% | 46.9 | 2465 | 35.9% |
| CoT-Valve | 92.70 | 526 | 9.9% | 84.2 | 2111 | 8.4% | 40.0 | 7027 | 29.0% | 41.7 | 2751 | 48.8% | 46.5 | 2872 | 25.3% |
| [1pt/2pt] (w/o D.H.) | 93.69 | 338 | 42.2% | 86.2 | 1549 | 32.8% | 46.7 | 7807 | 21.1% | 46.7 | 3745 | 30.3% | 50.5 | 2208 | 42.6% |
| (w/o R.H.) | 93.47 | 299 | 51.2% | 85.8 | 1706 | 26.0% | 50.0 | 8156 | 17.6% | 48.3 | 3858 | 28.2% | 50.5 | 2463 | 36.0% |
| TH2T | 94.08 | 246 | 58.1% | 86.4 | 1343 | 41.8% | 50.0 | 7140 | 27.9% | 48.3 | 3486 | 35.1% | 51.0 | 1901 | 50.6% |
| DeepSeek-R1-Distill-Qwen-32B | |||||||||||||||
| Original | 95.60 | 717 | 87.2 | 2357 | 60.0 | 9605 | 50.0 | 5540 | 52.0 | 4384 | |||||
| [1pt/2pt] D-Prompt | 95.60 | 663 | 7.6% | 86.8 | 2198 | 6.8% | 56.7 | 9445 | 1.7% | 48.3 | 5135 | 7.3% | 49.5 | 3877 | 11.6% |
| NoThinking | 93.86 | 228 | 68.2% | 80.2 | 659 | 72.0% | 53.3 | 3208 | 66.6% | 45.0 | 2146 | 61.3% | 50.5 | 1627 | 62.9% |
| TokenSkip | 96.54 | 423 | 41.0% | 79.8 | 1567 | 33.5% | 50.0 | 5856 | 39.0% | 43.3 | 3730 | 32.7% | 47.4 | 2874 | 34.4% |
| CoT-Valve | 96.10 | 781 | 8.9% | 85.3 | 2263 | 3.9% | 53.3 | 7997 | 16.7% | 43.3 | 4663 | 15.8% | 49.0 | 3550 | 19.0% |
| [1pt/2pt] (w/o D.H.) | 96.08 | 361 | 49.7% | 87.4 | 1863 | 20.9% | 56.7 | 7737 | 19.4% | 48.3 | 3432 | 38.1% | 51.5 | 2513 | 42.7% |
| (w/o R.H.) | 95.75 | 323 | 55.0% | 86.8 | 2114 | 10.3% | 60.0 | 8155 | 15.1% | 45.0 | 3644 | 34.2% | 51.0 | 2809 | 35.9% |
| TH2T | 97.04 | 263 | 63.3% | 88.0 | 1733 | 26.5% | 56.7 | 7004 | 27.1% | 48.3 | 2998 | 45.9% | 52.0 | 2367 | 46.0% |
| Bench | Difficulty lv. | Cognition Acc. | Confidence | ||
| Original | TH2T | Original | TH2T | ||
| GSM8K | all easy | 14.9% | 98.3% | 15.46 | 17.19 |
| MATH | lv.1-2 (easy) | 9.8% | 93.4% | 16.09 | 16.97 |
| lv.4-5 (hard) | 12.0% | 91.5% | 15.72 | 16.25 | |
| AIME2024 | all hard | 16.7% | 96.7% | 16.64 | 16.89 |
| OmniMath | lv.1-2 (easy) | 20.0% | 96.7% | 16.41 | 17.05 |
| lv.4-5 (hard) | 13.3% | 93.3% | 16.94 | 17.76 | |
5.2 Results
@@
TH2T achieves a significant reduction in computational overhead without compromising performance. The in-distribution performance on GSM8K and MATH is illustrated in Tab. 1. Key observations include: 1) TH2T exhibits minimal accuracy degradation; more precisely, most settings witness slight improvement. 2) TH2T achieves the best reduction in reasoning length among methods with stable accuracy—up to 74% on GSM8K and 38% on MATH with the 7B model, translating to over 5 and 2 latency speedups, respectively. 3) Greater gains are achieved on easier tasks, validating TH2T’s ability to adapt reasoning effort based on task difficulty. 4) These observations are consistent across 7B, 14B and 32B models. As shown in Fig. 1, other methods (e.g., NoThinking and CoT-Value) achieve comparable compression ratios in token length but at the cost of severe accuracy degradation (often 10 points), underscoring TH2T’s robustness in balanced efficiency.
TH2T substantially strengthens the calibration of task difficulty cognition. The entirety of GSM8K, MATH lv.1-2, and Omni-Math lv.1-2 are set as “Easy” while the remaining are “Hard”. As shown in Tab. 2 and Fig. 2, the original LRM reveals ambiguous recognition of task difficulty (typically ), defaulting to the Medium label in the 3-option setting and lacking discriminability in the 2-option setting. In contrast, our approach yields a clear distinction with precision above 90%, evidencing effective calibration of task difficulty cognition enabled by the first-stage training.
@@
TH2T substantially enhances cognition of redundancy. Tab. 3 reports the average number of reflective and looping structures, comparing R1-Distilled-7B and TH2T-7B. We observe that: 1) TH2T effectively suppresses the number of reflection chunks by 10 on easy tasks and 3 on hard tasks, suggesting that extensive reflection is superfluous for strong reasoning. 2) TH2T substantially alleviates the looping problem. Notably, 59.5% of incorrect answers by the original model on MATH contain tail recursion, which is reduced to 21.4% under TH2T. As shown in Fig. 5(b), we observe zero loop structure in GSM8K. This distinct pattern of short-length concentration highlights the efficacy of Stage-2 of TH2T, yielding more concise and efficient reasoning structures.
@@ 5.3 Analysis
Generalization Ability. To validate the generalization on OOD tasks, we evaluate on 2 more mathematical benchmarks, i.e., AIME2024 and OmniMath, and 1 QA benchmark, i.e., GPQA. As shown in Tab. 1, consistent with the main results, our approach preserves accuracy while yielding substantial token reduction up to 20-50%.
Internal states validate the enhanced cognition. To examine whether our method instills a genuine sense of difficulty cognition, we investigate the model’s internal state in Tab.2. We posit that the confidence Fu et al. (2025) of the first token acts as an intrinsic, metacognitive signal of the model’s perceived task difficulty. Our method exhibits higher confidence, aligning with a concurrent rise in difficulty assessment accuracy. This consistency serves as a clear, self-consistent internal signal, providing strong validation that the enhanced cognition is explicitly represented in internal state.
Ablation on two-stage design.
@@
We explore two more settings: 1) without Difficulty Hypnosis in Stage-1, and 2) without Redundancy Hypnosis in Stage-2, which are annotated as w/o D.H. and w/o R.H. in Tab. 1. We observe that D.H. introduces more token reduction on easy task GSM8K than on hard task MATH (e.g., 26.6% vs 4.8%, on 7B model), while R.H. exerts a greater impact on MATH (11.6% vs 17.3%). As we illustrate in Sec.4, we design Difficulty Cognition Injection Stage to build native perception on easy tasks and the Redundancy Cognition Injection Stage to mitigate non-essential structures that frequently occur in hard tasks. This phenomenon also holds for the other three datasets, demonstrating the rationality of our two-stage framework.
Acc and length across difficulty levels.
We report a breakdown of the accuracy and average token length distribution across the complete spectrum of difficulty levels on 4 tasks in Fig.6. We compare the original R1-Distilled-Qwen-7B with TH2T-7B, where higher numerical levels indicate greater question difficulty. Our observations are as follows: 1) Both models produce progressively longer and more elaborate responses with increasing difficulty. 2) Across all difficulty levels and tasks, TH2T consistently reduces response length while maintaining stable performance. 3) Length reduction is most pronounced for easier questions.
6 Conclusion